OpenStudy (anonymous):

Teas costing $7 per pound are to be mixed with teas costing $4 per pound to make a 30-pound mixture. If the mixture is to sell for $6 per pound, how many pounds of each kind of tea should be used?

OpenStudy (sriram):

let x pounds of $7 tea be used and y pounds of $4 pound. x+y=30 7x +4y=6*30=180

OpenStudy (sriram):

solving the eq simultaneously we get x=20 y=10

OpenStudy (sriram):

20 pound of $7 and 10 pound of $4

OpenStudy (anonymous):

thankss whats your name?

OpenStudy (sriram):

Sriram

OpenStudy (anonymous):

thank you :)